Sol Summary – February 20th

Crew 189 Sol Summary Report 20 Feb 2018

Sol 2

Summary Title: First experiments deployment

Author: Victoria Da-Poian

Mission Status: All nominal

Sol Activity Summary: Our day began with our lovely physical training.
It is already adopted by each crewmember. Since this morning, everyone
has experienced the EVAs. We managed to deploy one experiment (the
MegaARES) but a small problem enabled the LOAC deployment. We will fix
it with Jérémy in order to try again tomorrow. Even if Gabriel had a
very foggy helmet (fans problems), and everyone seems really happy and
I have to say the photos Benoit took are really nice. During our EVA,
the rest of the crew fixed the water boiler, cleaned, cooked and
worked on experiments. After the lunch and a little break, Gabriel
showed us his humans factors experiment, and everyone worked on their
own experiments and reports.

Look Ahead Plan: After our first deployment, the second air sensor,
(LOAC) will normally be deployed tomorrow.

Anomalies in work: Water leak in the Hab Lower deck, broken fans in spacesuits

Weather: Nice

Crew Physical Status: Fine, already a bit tired from the first EVAs

Reports to be filed
Greenhab Officer

Support Requested: No more boiler is needed. Would it be possible to
have the RoverCal that is in the RAM delivered in the airlock please?

Copyright © The Mars Society. All rights reserved. | Main Site